OpenStudy (anonymous):

Could someone help with manipulating and simplifying ths expression. a^2.b1/2.a^3.b2/3.a^0

OpenStudy (anonymous):

\[a^2 \times b \frac{ 1 }{ 2 } \times a^3 \times b \frac{ 2 }{ 3 } \times a ^0\]Is this your question?

jhonyy9 (jhonyy9):

how is it right,correct ? please us parantheses !!!

OpenStudy (anonymous):

Hi curiousshubham, yes that is my question could you explain how to simplify as i am new to this.

OpenStudy (anonymous):

Since multiplication is associative, you can rearrange the terms to put similar terms next to each other. You don't have to start that way, but it might make it easier to do the next step. \[a ^{2} \times b\frac{ 1 }{ 2 } \times a ^{3} \times b\frac{ 2 }{3} \times a ^{0}\] \[a ^{2} \times a ^{0} \times a ^{3} \times b\frac{ 1 }{ 2 } \times b\frac{ 2 }{3}\]

OpenStudy (anonymous):

Do you know the index rule:\[a^2 \times a^2 = a^{2+2} = a^4\]And next thing is: You can write: \[b \frac{ 1 }{ 2 } \rightarrow b \times \frac{ 1 }{ 2 }\] Then using the index rule, \[a^{2+3+0} \times b^{1+1} \times \frac{ 1 }{ 2 } \times \frac{ 2 }{ 3 }\]Now, ading up the power and simplifying the constants:\[\frac{ 1 }{ 3 } a^{5} b^{2}\]